OpenStudy (anonymous):

Help with functions if f(x^2-2x-4)=x-4, find f(x^2+6x+4)

OpenStudy (anonymous):

if f(x^2)=3x and f(-2x)=-2x and f(-4)=-4 then f(x^2-2x-4)=x-4, and f(x^2+6x+4)=3x +6x +4 = 9x+4